# The impact of Solar magnetic field configurations on the production of gamma rays at the Solar disk

### v1

@@ Appendix A Influence of the Magnetic Mirror in the Parker Field
In this section, we investigate the influence of the pitch-angle distribution (PAD) on the gamma-ray production. In the main part, we discussed an uniform PAD at the outer surface ($r=2.5R_{\odot}$). This assumption is valid in the case where the mean free path of the scattering $\lambda_{\mathrm{mfp}}$ is much shorter than the distance between Earth and Sun. Hutchinson-etal-2022 introduce the case of weak scattering, where the magnetic mirroring of GCRs within the Parker magnetic field becomes relevant. In the following, we introduce the PAD resulting from the mirroring in the Parker field (Sec. A.1) and estimate the changes in the gamma-ray flux based on the new PAD (Sec. A.2).

### v2

@@ Appendix A Influence of the Magnetic Mirror in the Parker Field
In this section, we investigate the influence of the pitch-angle distribution (PAD) on the gamma-ray production. In the main part, we discussed an uniform PAD at the outer surface ($r=2.5R_{\odot}$). This assumption is valid in the case where the mean free path of the scattering $\lambda_{\mathrm{mfp}}$ is much shorter than the distance between Earth and Sun. Hutchinson et al. (2022) introduce the case of weak scattering, where the magnetic mirroring of GCRs within the Parker magnetic field becomes relevant. In the following, we introduce the PAD resulting from the mirroring in the Parker field (Sec. A.1) and estimate the changes in the gamma-ray flux based on the new PAD (Sec. A.2).